# Scalable modular architecture for universal quantum computation

### v2

@@ IV Summary
Another option, even if information transfer between subsystems is slow, is to tailor the system architecture to run quantum algorithms that can be parallelized. Using each subsystem as a QPU, quantum circuits could be designed such that calculations are run in parallel on each subsystem and, at the end, all the information is merged before measuring. The parallelization of quantum algorithms is a current topic of discussion and has been hypothesized, for example, to allow for reduced circuit depths in some cases [43] or help with the implementation of decision diagrams [44].
In future work, it will also be interesting to extend the proof to systems that are not based on qubits. Throughout all lemmas, we have assumed two-level systems as elementary building blocks, allowing us to use tensor products of Pauli matrices as the basis for the dynamical Lie algebras. It thus needs to be shown that the same arguments can be made using e.g. the generalized skew-Hermitian Pauli matrices as a basis.
@@ Appendix A Proof of the theorem
To prove Theorem 1, we use an analytical approach to prove controllability of the multipartite system using the Lie rank condition [19]. This section proves that when adding any two-qubit tunable coupling between the subsystems $\mathcal{A}$ and $\mathcal{B}$, the whole system $\mathcal{A}\otimes\mathcal{B}$ is controllable.
@@ A.1 Notation and problem statement
Connecting two qubits from $\mathcal{A}$ and $\mathcal{B}$ by an entangling two-qubit coupling $w(t)\hat{H}^{\mu,n}_{c}$ (cf. Equation Eq. 2) that can be treated as an independent control, the Hamiltonian of the $\mathcal{A}\otimes\mathcal{B}$ system is given by
@@ A.3 Preliminaries
The next lemma includes a mention to the generation of a vector basis of a Lie algebra $\mathfrak{A}$ based on a set of elements $\{A_{1},A_{2},...A_{n}\}$. Here is a short description of one of the possible methods, extracted from described in Chapter 3 of [19]:
@@
Let $\{i\hat{A}_{\rho}\}_{\rho=0}^{C_{A}}\subset\mathfrak{su}(2^{M})$ be a set of linearly independent traceless skew-Hermitian operators that generate the algebra $\mathfrak{A}:=Lie\left[\{i\hat{A}_{\rho}\}_{\rho=0}^{C_{A}}\right]$. Let $i\hat{B}_{0}\in\mathfrak{su}(2^{N})$ be a nonzero operator such that $i\hat{A}_{\rho}\otimes\mathds{1}_{2^{N}}\,i\mathds{1}_{2^{M}}\otimes\hat{B}_{0}]=0$ for every $0\leq\rho\leq C_{A}$. Then the following implication holds true:
@@ A.4 Controllability of two controllable qubit arrays coupled via a two-qubit control
The Hamiltonians of systems $\mathcal{A}$ and $\mathcal{B}$ can be written as Equations (6) and (8), respectively. Thus the Hamiltonian of the bipartite system composed of subsystem $\mathcal{A}$, subsystem $\mathcal{B}$ and the tunable coupling $\hat{H}^{\mu,n}_{c}$ (cf. Equation Eq. 2) follows the form of Equation Section A.1.
@@
$\mathcal{L}_{\mathcal{A}\otimes\mathcal{B}}=$ $\;Lie\left[1,\;\{i\hat{A}_{\rho}\otimes\mathds{1}_{2^{N}}\}_{\rho=1}^{C_{A}},\;\{i\mathds{1}_{2^{M}}\otimes\hat{B}_{j}\}_{j=1}^{C_{B}},\;i\hat{H}^{\mu,n}_{c}\right]=$
@@
As a final note and without an explicit proof, this result can also be extended to any type of control that is entangling between the partitions $\mathcal{A}$ and $\mathcal{B}$. The main idea behind it is that the same process by which $i\hat{H}^{\mathcal{AB}}_{0}$ was reduced to a single product $c_{\alpha,j}\hat{\sigma}_{\alpha}^{(\mu)}\otimes\hat{\sigma}_{j}^{(n)}$ can be analogously defined for any type of entangling coupling.

### v3

@@ IV Summary and Perspectives
Another option, even if information transfer between subsystems is slow, is to tailor the system architecture to run quantum algorithms that can be parallelized. Using each subsystem as a QPU, quantum circuits could be designed such that calculations are run in parallel on each subsystem and, at the end, all the information is merged before measuring. The parallelization of quantum algorithms is a current topic of discussion and has been hypothesized, for example, to allow for reduced circuit depths in some cases [broadbent2009parallelizing] or help with the implementation of decision diagrams [li2024parallelizing].
In future work, it will be interesting to extend our proof to systems that are not based on qubits. In view of previous controllability results [zeier2011symmetry, albertini2025quantum], Theorem 1 holds also for quantum systems composed of subsystems with dimension other than $2^{N}$. Throughout all lemmas, we have assumed two-level systems as elementary building blocks, allowing us to use tensor products of Pauli matrices as the basis for the dynamical Lie algebras. The same arguments will then need to be made using a different basis of the Lie algebra, e.g. the set of generalized skew-Hermitian Pauli matrices. Finding a suitable set of operations such as the ones described in Appendix A.2 will allow for retrieving information on commutator depth for qudit-based systems.
A particularly relevant question for future work is a possible connection between the minimal commutator depth needed for controllability and the quantum speed limit of the system. Reducing the number of controls may lead to larger operation times, suggesting that there exists an optimal balance. In view of running quantum algorithms on actual devices, answering this question will also provide a link to the circuit depth of a quantum circuit transpiled to the native operations of the modular architecture.
Finally, it will be important to learn how errors propagate in a system when the connections between modules are reduced. Assuming that gates have a finite error, fewer connections imply more operations in a circuit, which may result in an overall lower fidelity. The exact impact of error propagation remains be explored in the future.
@@ Appendix A Proof of the theorem
To prove Theorem 1, we use an analytical approach to prove controllability of the multipartite system using the Lie rank condition [dAlessandro2021]. This section proves that when adding any two-qubit tunable coupling between the subsystems $\mathcal{A}$ and $\mathcal{B}$, the whole system $\mathcal{A}\otimes\mathcal{B}$ is controllable.
@@ A.1 Notation and problem statement
Connecting two qubits from $\mathcal{A}$ and $\mathcal{B}$ by an entangling two-qubit coupling $w(t)\hat{H}^{\mu,n}_{c}$ (cf. Eq. 2) that can be treated as an independent control, the Hamiltonian of the $\mathcal{A}\otimes\mathcal{B}$ system is given by
@@ A.3 Preliminaries
The next lemma includes a mention to the generation of a vector basis of a Lie algebra $\mathfrak{A}$ based on a set of elements $\{A_{1},A_{2},...A_{n}\}$. Here is a short description of one of the possible methods, described in Chapter 3 of [dAlessandro2021]:
@@
The following lemma shows that the drifts of the two subsystems used can always be decoupled. This is not straightforward, as the only control that we have over a drift is to let it evolve over time. When dealing with drifts of different subsystems, we can still prove that the two terms can always be decoupled in the algebra.
@@
Let $\{i\hat{A}_{\rho}\}_{\rho=0}^{C_{A}}\subset\mathfrak{su}(2^{M})$ be a set of linearly independent traceless skew-Hermitian operators that generate the algebra $\mathfrak{A}:=Lie\left[\{i\hat{A}_{\rho}\}_{\rho=0}^{C_{A}}\right]$. Let $i\hat{B}_{0}\in\mathfrak{su}(2^{N})$ be a nonzero operator such that $i\hat{A}_{\rho}\otimes\mathds{1}_{2^{N}}\,,\,i\mathds{1}_{2^{M}}\otimes\hat{B}_{0}]=0$ for every $0\leq\rho\leq C_{A}$. Then the following implication holds true:
@@ A.4 Controllability of two controllable qubit arrays coupled via a two-qubit control
The Hamiltonians of systems $\mathcal{A}$ and $\mathcal{B}$ can be written as Equations (6) and (8), respectively. Thus the Hamiltonian of the bipartite system composed of subsystem $\mathcal{A}$, subsystem $\mathcal{B}$ and the tunable coupling $\hat{H}^{\mu,n}_{c}$ (cf. Eq. 2) follows the form of Section A.1.
@@
$\mathcal{L}_{\mathcal{A}\otimes\mathcal{B}}=$ $\;Lie\left[\hat{H}_{0}^{AB},\;\{i\hat{A}_{\rho}\otimes\mathds{1}_{2^{N}}\}_{\rho=1}^{C_{A}},\;\{i\mathds{1}_{2^{M}}\otimes\hat{B}_{j}\}_{j=1}^{C_{B}},\;i\hat{H}^{\mu,n}_{c}\right]=$
@@
As a final note and without an explicit proof, this result can also be extended to any type of control that is entangling between the partitions $\mathcal{A}$ and $\mathcal{B}$. The main idea behind it is that the same process by which $i\hat{H}^{\mathcal{AB}}_{c}$ was reduced to a single product $c_{\alpha,j}\hat{\sigma}_{\alpha}^{(\mu)}\otimes\hat{\sigma}_{j}^{(n)}$ can be analogously defined for any type of entangling coupling.